OpenStudy (anonymous):

Finding the number of solutions to 0=8x^2

OpenStudy (anonymous):

Since the only solution is 0, the number of solutions is 1.

OpenStudy (anonymous):

it's not zero

OpenStudy (anonymous):

but how do you find it

OpenStudy (anonymous):

Divide each term in the equation by 8 8x^2/8=0/8 Simplify the left side of the equation by canceling the common factors x^2=0/8 simplify it x^2=0 Take the square root of both sides of the equation and eliminate the exponent on the left \[x=\pm \sqrt{0}\] \[x=\pm0\] \[\pm0=0\] x=0

OpenStudy (anonymous):

hope that helps

OpenStudy (anonymous):

zero is the number solution